# Spectral gap for the signed interchange process with arbitrary sets

### v2

@@ Introduction
Let $w=w_{N}+w_{T}$ be an element of the group ring $\mathbb{R}[B_{n}]$, where $w_{T}$ is a nonnegative combination of transpositions $(ij)\in S_{n}$, and $w_{N}$ is a nonnegative combination of the elements $s_{\{i\}}$, for $1\leq i\leq n$. Then we have
@@
Let $w=w_{N}+w_{T}$ where $w_{T}=\sum_{1\leq i<j\leq n}a_{ij}(ij)$ and $w_{N}=\sum_{A\subseteq\{1,\dotsc,n\}}\alpha_{A}s_{A}$ with any nonnegative weights $(a_{ij})$ and $(\alpha_{A})$. Then we have
$\psi_{B_{n}}(w)=\min\{\psi_{B_{n}}(w,\sigma)\,|\,\sigma\in\mathcal{F}_{n}\}.$
@@
In fact, we shall prove a slightly stronger result:
@@
Our final result shows that all of the irreps in $\mathcal{F}_{n}$ are required in Theorem 1, even when $w_{N}$ is only supported on odd sets:
@@
Let $\sigma\in\mathcal{F}_{n}$. Then there exists $w=w_{T}+w_{N}^{-}\in\mathbb{R}[B_{n}]$ where $w_{T}$ is a nonnegative linear combination of transpositions and $w_{N}^{-}$ is a nonnegative linear combination of $\{s_{A}\,|\,A\subseteq\{1,\dotsc,n\}\text{ and }|A|\text{ is odd }\}$ such that $\psi_{B_{n}}(w,\sigma)<\psi_{B_{n}}(w,\tau)$ for all nontrivial $B_{n}$–irreps $\tau\ncong\sigma$.
@@
The main part in the proof of Theorem 3 is the case where $\sigma=([n-k],[k])$ for some $k>0$. In that case we use Fourier theory on the hypercube to construct $w_{N}^{-}$ such that $\psi_{G}(w_{N}^{-},\sigma)<\psi_{G}(w_{N}^{-},([n-i],[i]))$ for all $0<i\neq k$, and then show that $w_{T}$ can be chosen as a certain multiple of the complete graph element $\sum_{1\leq i<j\leq n}(ij)$ to satisfy the required property.
@@ 1. Representations of $B_{n}$
Let $i=|\sigma_{1}|$, so $n-i=|\sigma_{2}|$. Pulling back by the quotient map $B_{i}\rightarrow B_{i}/N_{i}\cong S_{i}$, we consider $V_{\sigma_{1}}$ as a representation of $B_{i}$. Similarly, we consider $V_{\sigma_{2}}$ as a representation of $B_{n-i}$. Let $U$ be the $1$-dimensional representation of $B_{n-i}$, given on the $(n-i)\times(n-i)$ matrices by
$U(A)=\prod_{1\leq j,k\leq n-i\text{ : }A_{jk}\neq 0}A_{jk}.$
Then (Cf. [6], (3.1), and [10], Section 8.2), the $B_{n}$–irrep corresponding to the pair $(\sigma_{1},\sigma_{2})$ is
$V_{(\sigma_{1},\sigma_{2})}:=\operatorname{Ind}^{B_{n}}_{B_{i}\times B_{n-i}}((V_{\sigma_{1}}\boxtimes(V_{\sigma_{2}}\otimes U)))$ (2)
@@
    For any $\pi\in S_{n}$ and $A\subseteq\{1,\dotsc,n\}$ of size $i$, $\pi\cdot A=\pi(A)=\{\pi(i)\,|\,i\in A\}$.
@@
As $V_{[n-i]}$, $V_{[i]}$ and $U$ are one dimensional, so is the $B_{n-i}\times B_{i}$–irrep $V_{[n-i]}\boxtimes(V_{[i]}\otimes U)$, and the induction
$V_{([n-i],[i])}=\operatorname{Ind}_{B_{n-i}\times B_{i}}^{B_{n}}(V_{[n-i]}\boxtimes(V_{[i]}\otimes U))\cong\mathbb{R}[B_{n}]\underset{\mathbb{R}[B_{n-i}\times B_{i}]}{\otimes}(1\boxtimes U)$
@@
$=$ $\vartheta_{A}\cdot(1\boxtimes U)(s_{\vartheta_{A}^{-1}(B)})$
@@
$\pi\vartheta_{A}=\vartheta_{\pi(A)}\pi^{\prime}=\vartheta_{\pi(A)}(1\boxtimes U)(\pi^{\prime})=\vartheta_{\pi(A)}$
This completes the proof. ∎
@@ 2. Proof of Theorems 1 and 2
As a map between vector spaces, $\phi$ is nothing but the Fourier transform in $N_{n}$, therefore it is an isomorphism. Let us verify that this is also a map of $B_{n}$–representations. For any $A\subseteq\{1,\dots,n\}$, we have
@@
We now recall and prove Theorems 1 and 2:
Theorem 1.
Let $w=w_{N}+w_{T}$ where $w_{T}=\sum_{1\leq i<j\leq n}a_{ij}(ij)$ and $w_{N}=\sum_{A\subseteq\{1,\dotsc,n\}}\alpha_{A}s_{A}$ with any nonnegative weights $(a_{ij})$ and $(\alpha_{A})$. Then we have
$\psi_{B_{n}}(w)=\min\{\psi_{B_{n}}(w,\sigma)\,|\,\sigma\in\mathcal{F}_{n}\}.$
@@
where (1) follows since $\sigma(\Delta(w_{N}))$ is a nonnegative operator; (2) follows since $w_{T}$ is supported on $S_{n}$, and $\operatorname{Res}^{B_{n}}_{S_{n}}\sigma$ does not contain copies of the trivial $S_{n}$–irrep; (3) is immediate; (4) follows from the Aldous-Caputo-Liggett-Richthammer theorem; (5) is the same as (2) where $\sigma=([n-1,1],0)$; and (6) holds as the action matrix of $w_{N}$ on the representation $([n-1,1],\emptyset)$ is equal to 0. ∎
Finally, we note that Theorem 1 is an immediate consequence of Theorem 2.
@@ 3. Proof of Theorem 3
We now turn to showing by explicit construction that for any $n$, all the irreps in $\mathcal{F}_{n}$ are needed in order to account for the spectral gap of the general group ring elements of the form $w_{T}+w_{N}^{-}$.
@@
Let $\sigma\in\mathcal{F}_{n}$. Then there exists $w=w_{T}+w_{N}^{-}\in\mathbb{R}[B_{n}]$ where $w_{T}$ is a nonnegative linear combination of transpositions and $w_{N}^{-}$ is a nonnegative linear combination of $\{s_{A}\,|\,A\subseteq\{1,\dotsc,n\}\text{ and }|A|\text{ is odd }\}$ such that $\psi_{B_{n}}(w,\sigma)<\psi(w,\tau)$ for all nontrivial $B_{n}$–irreps $\tau\ncong\sigma$.
@@
Let us start with the main case $\sigma=([n-k],[k])$, where $1\leq k\leq n$. Let $A=\{1,2,\dotsc,k\}$ and $B=\{1,2,\dotsc n\}$. Consider the collection
@@
Let $w_{N}^{-}=4\sum_{S\in U}U$ and $w_{T}=\frac{2^{n}}{n}\sum_{1\leq i<j\leq n}(ij)$. We will show that $w=w_{T}+w_{N}^{-}$ has the desired properties.
@@
Next, we will find an upper bound for $\psi_{B_{n}}(w,([n-k],[k]))$. Let us consider the element
$v=\sum_{S\subseteq B,|S|=k}S\in V_{k}$
@@
Consider the standard inner product on $V_{n}^{k}$:
$\langle\sum\alpha_{S}S,\sum\beta_{S}S\rangle=\sum\alpha_{s}\beta_{s}$
$\psi_{B_{n}}(w,V_{n}^{k})=\min_{\xi\in V_{n}^{k}\setminus\{0\}}\frac{\langle\Delta(w)\xi,\xi\rangle}{\langle\xi,\xi\rangle}$
By (12) and (13),
@@
$=$ $\begin{cases}2^{n}\binom{n}{k}&\text{if }2k=n\\ (2^{n}-1)\binom{n}{k}&\text{otherwise}\end{cases}$
Hence, $\langle\Delta v,v\rangle\leq 2^{n}\binom{n}{k}$. As $\langle v,v\rangle=\binom{n}{k}$, we get
$\psi_{B_{n}}(w,V_{n}^{k})\leq\frac{\langle\Delta v,v\rangle}{\langle v,v\rangle}\leq 2^{n}$ (14)
@@
Together with (10), we get that $\psi_{B_{n}}(w,([n-k],[k]))$ is the unique minimum among the numbers $\psi_{B_{n}}(w,([n-i],[i]))$, for $1\leq i\leq n-1$.
@@
This concludes the case $\sigma=([n-k],[k])$.
@@
  • Case 1: $\tau_{2}\neq\emptyset$. In that case, by Lemma 8, $\operatorname{Res}^{B_{n}}_{N_{n}}\tau$ does not contain a copy of the trivial $N_{n}$–irrep. As $N_{n}$ is abelian, $\operatorname{Res}^{B_{n}}_{N_{n}}\tau$ factors into nontrivial one-dimensional irreps. By (3), these are all included in $\bigoplus_{i=1}^{n}V_{n}^{i}$. Therefore, using (16), we have
@@
    Let $\succeq$ be the dominance partial order on Young diagrams (a definition can be found, e.g., in [7] before Lemma 9). By Lemma 10 in [7], if $n\vdash\lambda,\lambda^{\prime}$ and $\lambda\succeq\lambda^{\prime}$ then $r(\lambda)\geq r(\lambda^{\prime})$. The proof of that lemma shows that in the case where $\lambda\neq\lambda^{\prime}$ we have a strict inequality $r(\lambda)>r(\lambda^{\prime})$.
    We have $n-1,1]\succ\tau_{1}$, and therefore, (18) holds.

### v4

@@ Introduction
Let $w=w_{T}+w_{N}$ be an element of the group ring $\mathbb{R}[B_{n}]$, where $w_{T}$ is a nonnegative combination of transpositions $(ij)\in S_{n}$, and $w_{N}$ is a nonnegative combination of the elements $s_{\{i\}}$, for $1\leq i\leq n$. Then we have
@@
and
$\mathcal{F}_{n}^{\raisebox{1.0pt}{$\scriptstyle-$}}=\{([n-1,1],\emptyset)\}\cup\{([n-k],[k])\,|\,1\leq k<n\}=\mathcal{F}_{n}\setminus\{(\emptyset,[n])\}$
@@
Let $w_{T}=\sum_{1\leq i<j\leq n}a_{ij}(ij)$ and $w_{N}=\sum_{A\subseteq\{1,\dotsc,n\}}\alpha_{A}s_{A}$ with any nonnegative weights $(a_{ij})$ and $(\alpha_{A})$. Let $w_{N}^{-}=\sum_{A\subseteq\{1,\dotsc,n\}\,:\,|A|\text{ is odd}}\alpha_{A}s_{A}$.
Then we have:
  1. (1)
    $\psi_{B_{n}}(w_{T}+w_{N})=\min\{\psi_{B_{n}}(w_{T}+w_{N},\sigma)\,|\,\sigma\in\mathcal{F}_{n}\}.$
  2. (2)
    $\psi_{B_{n}}(w_{T}+w_{N}^{-})=\min\{\psi_{B_{n}}(w_{T}+w_{N}^{-},\sigma)\,|\,\sigma\in\mathcal{F}_{n}^{\raisebox{1.0pt}{$\scriptstyle-$}}\}.$
@@
Our proof of Theorem 1 relies on the following result, which may be of independent interest:
@@
Our next result shows that all of the irreps in $\mathcal{F}_{n}$ (and respectively, $\mathcal{F}_{n}^{\raisebox{1.0pt}{$\scriptstyle-$}}$) are required in Theorem 1:
@@
  1. (1)
    Let $\sigma\in\mathcal{F}_{n}$. Then there exist $w_{T},w_{N}$ as in Theorem 1 such that $\tau=\sigma$ is the unique minimizer of the expression $\psi_{B_{n}}(w_{T}+w_{N},\tau)$ among all nontrivial irreps $\tau$ of $B_{n}$.
  2. (2)
    Let $\sigma\in\mathcal{F}_{n}^{\raisebox{1.0pt}{$\scriptstyle-$}}$. Then there exist $w_{T},w_{N}^{-}$ as in Theorem 1 such that $\tau=\sigma$ is the unique minimizer of the expression $\psi_{B_{n}}(w_{T}+w_{N}^{-},\tau)$ among all nontrivial irreps $\tau$ of $B_{n}$.
@@
The main part in the proof of Theorem 3 is the case where $\sigma=([n-k],[k])$ for some $1\leq k\leq n-1$. In that case we use Fourier theory on the hypercube to construct $w_{N}^{-}$ such that $\psi_{G}(w_{N}^{-},\sigma)<\psi_{G}(w_{N}^{-},([n-i],[i]))$ for all $0<i\neq k$, and then show that $w_{T}$ can be chosen as a certain multiple of the complete graph element $\sum_{1\leq i<j\leq n}(ij)$ to satisfy the required property.
@@ 1. Representations of $B_{n}$
We start by defining a multiplicative character $\nu_{n}$ of $B_{n}$: It is given on the $n\times n$ matrices by
$\nu_{n}(A)=\prod_{1\leq i,j\leq n\text{ : }A_{ij}\neq 0}A_{ij}.$ (2)
@@
Let $i=|\sigma_{1}|$, so $n-i=|\sigma_{2}|$. Pulling back by the quotient map $B_{i}\rightarrow B_{i}/N_{i}\cong S_{i}$, we consider $V_{\sigma_{1}}$ as a representation of $B_{i}$. Similarly, we consider $V_{\sigma_{2}}$ as a representation of $B_{n-i}$. Then (Cf. [6], (3.1), and [10], Section 8.2), the $B_{n}$–irrep corresponding to the pair $(\sigma_{1},\sigma_{2})$ is
$V_{(\sigma_{1},\sigma_{2})}:=\operatorname{Ind}^{B_{n}}_{B_{i}\times B_{n-i}}((V_{\sigma_{1}}\boxtimes(V_{\sigma_{2}}\otimes\nu_{n-i})))$ (3)
@@
    For any $\pi\in S_{n}$ and $A\subseteq\{1,\dotsc,n\}$ of size $i$, $\pi\cdot A=\pi(A)=\{\pi(j)\,|\,j\in A\}$.
@@
Consider the trivial representations $V_{[n-i]}$, $V_{[i]}$ of $S_{n-i}$ and $S_{i}$, respectively. Pulling back through the quotient maps $B_{n-i}\rightarrow S_{n-i}$ and $B_{i}\rightarrow S_{i}$, we view them as the trivial representations of $B_{n-i}$ and $B_{i}$, respectively. As $V_{[n-i]}$, $V_{[i]}$ and $\nu_{i}$ are one dimensional, so is the $B_{n-i}\times B_{i}$–irrep $V_{[n-i]}\boxtimes(V_{[i]}\otimes\nu_{i})$, and the induction
$V_{([n-i],[i])}=\operatorname{Ind}_{B_{n-i}\times B_{i}}^{B_{n}}(V_{[n-i]}\boxtimes(V_{[i]}\otimes\nu_{i}))\cong\mathbb{R}[B_{n}]\underset{\mathbb{R}[B_{n-i}\times B_{i}]}{\otimes}(1\boxtimes\nu_{i})$
@@
$=$ $\vartheta_{A}\cdot(1\boxtimes\nu_{i})(s_{\vartheta_{A}^{-1}(B)})$
@@
$\pi\vartheta_{A}=\vartheta_{\pi(A)}\pi^{\prime}=\vartheta_{\pi(A)}(1\boxtimes\nu_{i})(\pi^{\prime})=\vartheta_{\pi(A)}$
Therefore, the action of the permutations and the diagonal elements $s_{A}$ on the coset representatives $\vartheta_{A}$ is compatible with their action on the sets of cardinality $i$ in definition 5. This completes the proof. ∎
@@ 2. Proof of Theorems 1 and 2
As a map between vector spaces, $\phi$ is nothing but the Fourier transform in $N_{n}$, therefore it is an isomorphism. Let us verify that this is also a map of $B_{n}$–representations. We shall denote by $\oplus$ the symmetric difference operation between sets. For any $A\subseteq\{1,\dots,n\}$, we have
@@
We now recall and prove Theorems 2 and 1:
@@
where (1) follows since $\sigma(\Delta(w_{N}))$ is a nonnegative operator; (2) follows since $w_{T}$ is supported on $S_{n}$, and $\operatorname{Res}^{B_{n}}_{S_{n}}\sigma$ does not contain copies of the trivial $S_{n}$–irrep; (3) is immediate; (4) follows from the Aldous-Caputo-Liggett-Richthammer theorem; (5) is the same as (2) where $\sigma=([n-1,1],\emptyset)$ (and also using Lemma 4); and (6) holds as the action matrix of $\Delta(w_{N})$ on the representation $([n-1,1],\emptyset)$ is equal to 0. ∎
Theorem 1.
Let $w_{T}=\sum_{1\leq i<j\leq n}a_{ij}(ij)$ and $w_{N}=\sum_{A\subseteq\{1,\dotsc,n\}}\alpha_{A}s_{A}$ with any nonnegative weights $(a_{ij})$ and $(\alpha_{A})$. Let $w_{N}^{-}=\sum_{A\subseteq\{1,\dotsc,n\}\,:\,|A|\text{ is odd}}\alpha_{A}s_{A}$.
Then we have:
  1. (1)
    $\psi_{B_{n}}(w_{T}+w_{N})=\min\{\psi_{B_{n}}(w_{T}+w_{N},\sigma)\,|\,\sigma\in\mathcal{F}_{n}\}.$
  2. (2)
    $\psi_{B_{n}}(w_{T}+w_{N}^{-})=\min\{\psi_{B_{n}}(w_{T}+w_{N}^{-},\sigma)\,|\,\sigma\in\mathcal{F}_{n}^{\raisebox{1.0pt}{$\scriptstyle-$}}\}.$
Proof.
The first part follows directly from Theorem 2. For the second part, let $w=w_{T}+w_{N}^{-}$. Let us assume, without loss of generality, that $\alpha_{A}=0$ for any $A$ of even size, so $w_{N}=w_{N}^{-}$.
By the first part, we have $\psi_{B_{n}}(w)=\min\{\psi_{B_{n}}(w,\sigma)\,|\,\sigma\in\mathcal{F}_{n}\}.$ Let $\sigma_{0}=(\emptyset,[n])$. Since $\mathcal{F}_{n}^{\raisebox{1.0pt}{$\scriptstyle-$}}=\mathcal{F}_{n}\setminus\{\sigma_{0}\}$, it suffices to show that there exists $\tau\in\mathcal{F}_{n}^{\raisebox{1.0pt}{$\scriptstyle-$}}$ such that $\psi_{B_{n}}(w,\tau)\leq\psi_{B_{n}}(w,\sigma_{0})$. In fact, we shall now prove that this holds for any $\tau=([n-k],[k])$ with $1\leq k\leq n-1$.
By (3), $\sigma_{0}$ is the one-dimensional representation given by $\nu_{n}$ (see also (2)). Since $\nu_{n}(s_{A})=-1$ for any $A$ of odd size, and $\nu_{n}(\pi)=1$ for any $\pi\in S_{n}$, we have
$\psi_{B_{n}}(w,\sigma_{0})=\nu_{n}(\Delta(w))=\sum_{A\subseteq\{1,\dotsc,n\}}\alpha_{A}(1-\nu_{n}(s_{A}))=2\sum_{A\subseteq\{1,\dotsc,n\}}\alpha_{A}$ (5)
By Lemma 6, $\tau$ is isomorphic to the irrep $V_{n}^{k}$. Consider the standard inner product on $V_{n}^{k}$:
$\langle\sum\alpha_{S}S,\sum\beta_{S}S\rangle=\sum\alpha_{s}\beta_{s}$ (6)
$\psi_{B_{n}}(w,V_{n}^{k})=\min_{\xi\in V_{n}^{k}\setminus\{0\}}\frac{\langle\Delta(w)\xi,\xi\rangle}{\langle\xi,\xi\rangle}$ (7)
Let us consider the element
$v=\sum_{S\subseteq\{1,\dotsc,n\},|S|=k}S\in V_{n}^{k}$
As $v$ is $S_{n}$-invariant, we have $\Delta(w_{T})v=0$. For any $A\subseteq\{1,\dotsc,n\}$ we have by (4),
$s_{A}v=\sum_{S\subseteq\{1,\dotsc,n\},|S|=k}(-1)^{|A\cap S|}S$
and therefore,
$\langle s_{A}v,v\rangle=\sum_{S\subseteq\{1,\dotsc,n\},|S|=k}(-1)^{|A\cap S|}\geq\sum_{S\subseteq\{1,\dotsc,n\},|S|=k}(-1)=-\binom{n}{k}$.
We also have
$\langle v,v\rangle=\binom{n}{k}$ (8)
Therefore:
$\langle\Delta(w)v,v\rangle$ $=$ $\langle\Delta(w_{T})v,v\rangle+\langle\Delta(w_{N}^{-})v,v\rangle$
$=$ $\langle\Delta(w_{N}^{-})v,v\rangle=\sum_{A\subseteq\{1,\dotsc,n\}}\alpha_{A}\langle v-s_{A}v,v\rangle$
$=$ $\sum_{A\subseteq\{1,\dotsc,n\}}\alpha_{A}(\langle v,v\rangle-\langle s_{A}v,v\rangle)\leq 2\binom{n}{k}\sum_{A}\alpha_{A}$
By (5), (7) and (8), we conclude:
$\psi_{B_{n}}(w,([n-k],[k]))\leq\frac{\langle\Delta(w)v,v\rangle}{\langle v,v\rangle}\leq\frac{2\binom{n}{k}\sum_{A}\alpha_{A}}{\binom{n}{k}}=\psi_{B_{n}}(w,\sigma_{0}).$
@@ 3. Proof of Theorem 3
We now turn to showing by explicit construction that for any $n$, all the irreps in $\mathcal{F}_{n}$ (resp. $\mathcal{F}_{n}^{\raisebox{1.0pt}{$\scriptstyle-$}}$) are needed in order to account for the spectral gap of the general group ring elements of the form $w_{T}+w_{N}$ (resp. $w_{T}+w_{N}^{-}$).
@@
  1. (1)
    Let $\sigma\in\mathcal{F}_{n}$. Then there exist $w_{T},w_{N}$ as in Theorem 1 such that $\tau=\sigma$ is the unique minimizer of the expression $\psi_{B_{n}}(w_{T}+w_{N},\tau)$ among all nontrivial irreps $\tau$ of $B_{n}$.
  2. (2)
    Let $\sigma\in\mathcal{F}_{n}^{\raisebox{1.0pt}{$\scriptstyle-$}}$. Then there exist $w_{T},w_{N}^{-}$ as in Theorem 1 such that $\tau=\sigma$ is the unique minimizer of the expression $\psi_{B_{n}}(w_{T}+w_{N}^{-},\tau)$ among all nontrivial irreps $\tau$ of $B_{n}$.
@@
Let us start with the second part of the theorem, and with the main case $\sigma=([n-k],[k])$, where $1\leq k\leq n-1$. Let $A=\{1,2,\dotsc,k\}$ and $B=\{1,2,\dotsc n\}$. Consider the collection
@@
Let $w_{N}^{-}=4\sum_{C\in U}s_{C}$ and $w_{T}=\frac{2^{n}}{n}\sum_{1\leq i<j\leq n}(ij)$. We will show that $w=w_{T}+w_{N}^{-}$ satisfies the desired property, namely, that $\psi_{B_{n}}(w,\sigma)<\psi_{B_{n}}(w,\tau)$ for any nontrivial $B_{n}$–irrep $\tau\ncong\sigma$.
@@
Next, we will find an upper bound for $\psi_{B_{n}}(w,([n-k],[k]))$. Let us consider (as in the proof of Theorem 1) the element
$v=\sum_{S\subseteq B,|S|=k}S\in V_{n}^{k}$
@@
Recall the standard inner product on $V_{n}^{k}$ ((6)). By (17) and (18),
@@
$=$ $\begin{cases}2^{n}\binom{n}{k}&\text{if }2k=n\\ 2^{n}(\binom{n}{k}-1)&\text{otherwise}\end{cases}$
Hence, $\langle\Delta(w)v,v\rangle\leq 2^{n}\binom{n}{k}$. As $\langle v,v\rangle=\binom{n}{k}$, we get by the Rayleigh quotient principle ((7)):
$\psi_{B_{n}}(w,V_{n}^{k})\leq\frac{\langle\Delta(w)v,v\rangle}{\langle v,v\rangle}\leq 2^{n}$ (19)
@@
Together with (15), we get that $\psi_{B_{n}}(w,([n-k],[k]))$ is the unique minimum among the numbers $\psi_{B_{n}}(w,([n-i],[i]))$, for $1\leq i\leq n$.
@@
This concludes the case $\sigma=([n-k],[k])$ in the second part of the theorem.
@@
  • Case 1: $\tau_{2}\neq\emptyset$. In that case, by Lemma 8, $\operatorname{Res}^{B_{n}}_{N_{n}}\tau$ does not contain a copy of the trivial $N_{n}$–irrep. As $N_{n}$ is abelian, $\operatorname{Res}^{B_{n}}_{N_{n}}\tau$ decomposes into nontrivial one-dimensional irreps. By (4), these are all included in $\bigoplus_{i=1}^{n}V_{n}^{i}$. Therefore, using (21), we have
@@
    Let $\succeq$ be the dominance partial order on Young diagrams (a definition can be found, e.g., in [7] before Lemma 9). By Lemma 10 in [7], if $n\vdash\lambda,\delta$ and $\lambda\succeq\delta$ then $r(\lambda)\geq r(\delta)$. The proof of that lemma shows that in the case where $\lambda\neq\delta$ we have a strict inequality $r(\lambda)>r(\delta)$.
    We have $n-1,1]\succ\tau_{1}$, and therefore, (23) holds. This concludes the proof of the second part of the theorem.
As for the first part of the theorem, let $\sigma\in\mathcal{F}_{n}$. If $\sigma\in\mathcal{F}_{n}^{\raisebox{1.0pt}{$\scriptstyle-$}}$ then we are done by the second part of the theorem. Therefore, it is enough to prove the claim for $\sigma=\sigma_{0}=(\emptyset,[n])$. For that matter, let us define $w=w_{T}+w_{N}$ where $w_{T}=\sum_{1\leq i<j\leq n}(ij)$ and $w_{N}=\sum_{B\subseteq\{1,\dotsc,n\},\,|B|=2}s_{B}$. As we have seen in the proof of Theorem 1,
$\psi_{B_{n}}(w,\sigma_{0})=\nu_{n}(w)=0.$
Therefore, it suffices to show that $\psi_{B_{n}}(w,\tau)>0$ for any nontrivial $B_{n}$-irrep $\tau\ncong\sigma_{0}$. Let us consider the following cases:
  • Case 1: $\tau=([n-k],[k])$ for some $1\leq k\leq n-1$. For each $A\subseteq\{1,\dotsc,n\}$ of size $k$, and $B\subseteq\{1,\dotsc,n\}$ of size $2$, we have
    $\Delta(s_{B})A=(1-(-1)^{|A\cap B|})A=\begin{cases}2A&\text{if }|A\cap B|=1\\ 0&\text{otherwise}\end{cases}$
    For each such $A$ there are k(nk)k(n-k) subsets $B$ such that $|A\cap B|=1$. Therefore,
    $\Delta(w_{N})A=\sum_{B\subseteq\{1,\dotsc,n\},\,|B|=2}\Delta(s_{B})A=2k(n-k)A.$
    We conclude that $\Delta(w_{N})$ is a scalar operator on $V_{n}^{k}$ with a single eigenvalue 2k(nk)2k(n-k). Therefore,
    $\psi(w,\tau)\geq\psi(w_{N},\tau)=2k(n-k)>0.$
  • Case 2: $\tau=([n-1,1],\emptyset)$. By Lemmas 4 and 9,
    $\psi_{B_{n}}(w,\tau)\geq\psi_{B_{n}}(w_{T},\tau)=\psi_{S_{n}}(w_{T},[n-1,1])=n>0.$
  • Case 3: $\tau\notin\mathcal{F}_{n}^{\raisebox{1.0pt}{$\scriptstyle-$}}$. We have assumed that $\tau\ncong\sigma_{0}$, therefore $\tau\notin\mathcal{F}_{n}$. By Theorem 2 and Case 2:
    $\psi_{B_{n}}(w,\tau)\geq\psi_{B_{n}}(w,([n-1,1],\emptyset))>0.$
As all the cases were covered, our proof is complete. ∎